In a class test, the sum of Shefali's marks in Mathematics and English is 30. Had she got 2 marks more in Mathematics and 3 marks less in English, the product of their marks would have been 210. Find her marks in the two subjects.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Let us say, the marks of Shefali in Maths bex.

Then, the marks in English will be 30 -x.

As per the given question,

(x+ 2)(30 -x- 3) = 210

(x+ 2)(27 -x) = 210

⇒-x^2+ 25x+ 54 = 210

⇒x^2- 25x+ 156 = 0

⇒x^2-12x-13x+156=0

⇒x(x- 12) -13(x- 12) = 0

⇒(x- 12)(x- 13) = 0

⇒x= 12, 13

Therefore, if the marks in Maths are 12, then marks in English will be 30 - 12 = 18,andthe marks inMaths are 13, then marks in English will be 30 - 13 = 17

Thanks


.